OpenStudy (anonymous):

Simplify the expressions below. Write the final product in standard form and show your work to receive full credit. (x2 + 4x – 3)(2x2 + x + 6) My answer for this is: 2x^4 + 9x^3 + x^2 + 25x - 18 but I don't think that's right, can someone help me figure out what I did wrong??

OpenStudy (anonymous):

Here's how I got my answer... (x^2)(2x^2) = 2x^4 (x^2)(x) = x^3 (x^2)(6) = 6x^2 (4x)(2x^2) = 8x^3 (4x)(x) = 4x (4x)(6) = 24x (-3)(2x^2) = -6x^2 (-1)(x) = -3x (-1)(6) = -18

OpenStudy (anonymous):

the length of the cuboid is 8cm the width is 5cm the height is hcm work out the value of h

OpenStudy (anonymous):